A 10Ω resistor is connected in parallel to another resistor R. The equivalent resistance of the pair is 8Ω. What is the resistance R?

a. 10⁢Ω

b. 20Ω

c. 30Ω

d. 40Ω
Answer #1

When two resistors R1 and R2 are connected in parallel, their equivalent resistance will be

R = (1/R1 + 1/R2 ​​​​​​)-1​​​​​​

Here 8 = (1/10 + 1/R2​​​​​)-1

1/8 = 1/10 + 1/R2

​1/R2 = 1/8 - 1/10

R2 = 10*8/ (10-8)

R2 = 40 ohm

Therefore the another resistor is 40 ohm